How much will $600 per year be worth in eight years at interest rates of

a.13%. Round FVFA values in intermediate calculations to four decimal places. Round the answer to 2 decimal places.

b.$ 7%. Round FVFA values in intermediate calculations to four decimal places. Round the answer to 2 decimal places.

c. $ 5%. Round FVFA values in intermediate calculations to four decimal places. Round the answer to 2 decimal places.
Answer #1

a.Information provided:

Annuity= $600

Time= 8 years

Interest rate= 13%

The future value of the annuity is calculated by entering the below in a financial calculator:

PMT= 600

N= 8

I/Y= 13

Press the CPT key and FV to compute the future value of the annuity.

The value obtained is 4,464.2248.

Therefore, the future value of the annuity is $4,464.22.

b.Annuity= $600

Time= 8 years

Interest rate= 7%

The future value of the annuity is calculated by entering the below in a financial calculator:

PMT= 600

N= 8

I/Y= 7

Press the CPT key and FV to compute the future value of the annuity.

The value obtained is 6,155.8815

Therefore, the future value of the annuity is $6,155.88.

c. Annuity= $600

Time= 8 years

Interest rate= 5%

The future value of the annuity is calculated by entering the below in a financial calculator:

PMT= 600

N= 8

I/Y= 5

Press the CPT key and FV to compute the future value of the annuity.

The value obtained is 5,729.4653.

Therefore, the future value of the annuity is $5,729.47.

In case of any query, kindly comment on the solution.